OpenStudy (anonymous):

what is the mass (in grams) of 9.31 x 10^24 molecules of methanol?

OpenStudy (anonymous):

convert to moles and then to grams

OpenStudy (anonymous):

Remember what I said Grams-Moles-Atoms

OpenStudy (anonymous):

So you will start at Atoms, then get Moles, than Gram

OpenStudy (anonymous):

*CH4

OpenStudy (anonymous):

CH4 mass= 32.04amu C=12.01,H=1.004*4 1.) 9.3*10^24CH4atoms * 1mole 6.022*10^23atoms =15.4434molesCH4 2.) 15.4434molesCH4*32.04 grams 1mole = 494.81gramsCH4

OpenStudy (anonymous):

Grams-moles-Atoms

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

!